1.     Suppose we have the utility function, U = XY + X + Y.

a.      Find the function for the marginal rate of substitution.

  1. If prices are Px = $2 and Py = $4, and if income is M = $18, find the utility maximizing consumption bundle.
1
Expert's answer
2021-11-18T10:24:14-0500

Solution:

1.). a.). MRS = "\frac{MU_{x} }{MU_{y} }"

MUx = "\frac{\partial U} {\partial X}" = Y + 1


MUy = "\frac{\partial U} {\partial Y}" = X + 1

 

MRS = "\frac{MU_{x} }{MU_{y} }" = "\frac{Y+1}{X+1 }"


MRS Function = "\frac{Y+1}{X+1 }"

 

1.). Budget constraint: I = PxX + PyY

18 = 2X + 4Y

Set MRS = Px"\div" Py

"\frac{Y+1}{X+1 } = \frac{2}{4}"


Y = "\frac{X}{2 } - \frac{1}{2}"


Substitute in the budget constraint to get X:

18 = 2X + 4Y

18 = 2X + 4("\frac{X}{2 } - \frac{1}{2}" )

X = 5


Y = "\frac{X}{2 } - \frac{1}{2} = \frac{5}{2 } - \frac{1}{2} = 2.5 - 0.5 = 2"


Y = 2

The utility-maximizing consumption bundle (UX,Y) = (5, 2)